One number is 14 more than another. When half of the smaller number is added to the larger number, the sum is 59. What are the numbers?
Answer by CubeyThePenguin(3113)

x = larger number
y = smaller number

x = y + 14
x + y/2 = 59

Use substitution.


y + 14 + y/2 = 59
3y/2 = 45
y = 30

The numbers are 44 and 30.